OpenStudy (anonymous):

which of the following are solutions to the equation below ? Check all that apply. 3x^2+27x=30 A. -2 B. 10 C. -1 D. -10 E. -2 F. 1

OpenStudy (anonymous):

HERE

OpenStudy (anonymous):

inbox

OpenStudy (anonymous):

is it A nd B

OpenStudy (anonymous):

cool thnxs

OpenStudy (anonymous):

so not A and B

OpenStudy (radar):

Factor getting (x + 10)(3x-3)= 3(x=10)(x-1) x=-10 x=1

OpenStudy (mathmate):

@romanortiz65 I suggest you stop making guesses. With six choices, knowing that there are two solutions, it will take you 15 tries, assuming someone will tell you which one is right.